40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Feb 7 15:46:42 2013, Jenne, Update, Locking, PR2 flipped 
    Reply  Fri Feb 8 02:12:14 2013, Jenne, Update, Locking, PRMI work 
       Reply  Fri Feb 8 02:38:04 2013, Koji, Update, Locking, PRMI work 
       Reply  Fri Feb 8 11:01:18 2013, Jamie, Update, Locking, PRMI work 
          Reply  Fri Feb 8 12:39:32 2013, yuta, Update, Locking, PRMI work 
          Reply  Fri Feb 8 15:53:48 2013, Jenne, Update, Locking, PRMI work 
             Reply  Fri Feb 8 17:41:34 2013, Jenne, Update, Locking, PRMI work 
                Reply  Fri Feb 8 19:39:02 2013, Koji, Update, Locking, PRMI work 
                Reply  Fri Feb 8 20:05:15 2013, Jenne, Update, Locking, PRMI work 
Message ID: 8032     Entry time: Fri Feb 8 11:01:18 2013     In reply to: 8030     Reply to this: 8034   8037
Author: Jamie 
Type: Update 
Category: Locking 
Subject: PRMI work 

I completely agree with Koji.  We definitely should have locked the half PRC first.  We were all set up for that.  Why go through all this work to align MICH when we haven't confirmed with the half PRC that the flipping is helping us?  The first rule of debugging is to only make one change at a time.  We have measurements from the half PRC, so we could have made a direct comparison with those to see how things have changed.  If we jump the gun we're going to end up wasting more time when we have to back-track.

Also, we never talked about moving PR2 to adjust optical path length, although I can understand why we would think that should be done.  My calculations were all done assuming the free-space separation between PRM/PR2 and PR2/PR3 were unchanged.  It's possible changing the position is better, but again, it's more work and it changes multiple things at one.  I can redo my calculations for this new scenario, but we need to update our drawings with this new configuration.  Please note precisely where PR2 has moved to.

We should have just flipped PR2 and that's it.  Then we could have run the exact same measurements we had previously.  Only then, once we understood this new simple cavity, should we have done further adjustments.

ELOG V3.1.3-